# SurFhead: Affine Rig Blending for Geometrically Accurate 2D Gaussian Surfel Head Avatars

### v1

@@ 2.1.1 Gaussian Splatting
3D Gaussian Splatting (Kerbl et al., 2023) (3DGS) reconstructs a scene with anisotropic 3D Gaussian primitives. Each Gaussian is defined by a positive semi-definite covariance matrix $\Sigma$ that is centered at a position $\mu$. The covariance $\Sigma$ is decomposed as $\Sigma=RSS^{T}R^{T}$, where $R$ is a rotation matrix and $S$ is a diagonal scaling matrix. This covariance matrix is used when estimating each Gaussians’s function value at 3D point $x$ such as $G(x)=\text{exp}({-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)})$. Besides these geometrical properties, each Gaussian has appearance properties opacity $\alpha$ and color $c$. To render the scene, the final color $C$ is computed by alpha-blending Gaussians after projecting them to the image plane: $C=\sum_{i=1}c_{i}\alpha_{i}G^{proj}(x)\prod_{j=1}^{i-1}(1-\alpha_{i}G^{proj}(x)).$ $G^{proj}$ is given by evaluating the function value of 2D projection of the 3D Gaussian in image space by EWA volume splatting algorithm (Zwicker et al., 2001).
@@ 2.2 Affine transformation of 2D surfels
Local Geometry Descriptor. Let us consider a triangle in the canonical space as the matrix formed by its vertices ${v_{i}}\in\mathbb{R}^{3}$, and its edge matrix as $\mathbf{E}=[v_{1}-v_{0},v_{2}-v_{0},v_{3}-v_{0}]\in\mathbb{R}^{3\times 3}$, where $v_{3}:=v_{0}+\frac{(v_{1}-v_{0})\times(v_{2}-v_{0})}{||(v_{1}-v_{0})\times(v_{2}-v_{0})||_{2}}$. $\tilde{\mathbf{E}}$ is the deformed version with the deformed triangle’s vertices $\tilde{v_{i}}$ following Eq. 1. Note that in GaussianAvatars, the edge matrix is defined with a normalized base, height, and normal direction. This constructs an orthogonal matrix and approximates the shear or stretching-related deformations with isotropic scaling from the average of the base and height length. Please find a detailed derivation in GaussianAvatars (Qian et al., 2024).
@@ 2.3 Jacobian Blend Skinning (JBS)
Only with the local affine transformations might introduce discontinuities among adjacent triangles. This subsection introduces an improved technique to alleviate such discontinuities: Jacobian Blend Skinning. Specifically, we blend the local deformation with adjacent triangles’ deformations as done in previous arts (Zielonka et al., 2023b; Shao et al., 2024).
@@
Degenerate Solution of Linear Blend Skinning in Matrix Space. Linear Blend Skinning (LBS) (Badler & Morris, 1982) is widely used for dynamic modeling of the human body and head (Li et al., 2017; Loper et al., 2023), but it struggles with the non-linearity of rotation matrix interpolation in $SO(3)$, causing distortions in rotational transformations. Our toy experiments (Fig. 2(b)) show how element-wise linear interpolation distorts shapes. Previous works (Zielonka et al., 2023b; Shao et al., 2024) also adopt this sub-optimal approach to handle local deformations. Techniques like matrix-logarithm interpolation or quaternion’s SLERP can address this but are limited to rotations.
@@
Introducing Jacobian Blend Skinning (JBS). The Jacobian Blend Skinning (JBS) algorithm overcomes the limitations of LBS by focusing on the interpolation of Jacobian gradients, which encapsulate not only rotations but also shear and stretch, both residing in the broader $GL(3)$ (General Linear) field. The key to JBS is leveraging Polar Decomposition (PD) (Shoemake & Duff, 1992) to break down transformations into two meaningful components: rotation (orthogonal matrix $\mathbf{U}$) and stretch/shear (symmetric positive semi-definite matrix $\mathbf{P}$). This decomposition is unique and coordinate-independent (proof in Appendix A.2.1), making it geometrically sound for interpolation. By blending the rotation in matrix-logarithm space and the stretch/shear in linear matrix space, JBS ensures that the resulting transformations remain geometrically valid, avoiding the distortions seen with LBS (Fig. 2(b)).
JBS is applied to the Gaussian deformations. As aforementioned in Eq. 1, each Gaussian has its own parents’ deformation parameters which correspond to the current expression and pose, such as isotropic scale $s_{p}$ and rotation $R_{p}$. These parameters compose the deformed square root covariance and position of Gaussian, $\Sigma^{1/2}=s_{p}R_{p}R_{c}S_{c}$ and $\mu=s_{p}R_{p}\mu_{c}+T_{p}$, each. We replace the deformation of covariance term $s_{p}R_{p}$ to the blended Jacobian, $\mathbf{J}_{b}$, which is computed by blending adjacent triangles’ deformations.
@@
The blending itself is performed separately for the rotational ($\mathbf{U}_{b}$) and stretch/shear ($\mathbf{P}_{b}$) components. For rotations, the blending weights are applied in the matrix-logarithm space $\mathfrak{so}(3)$, followed by an exponential mapping back into the $SO(3)$ space. For stretch/shear, the weights are applied directly in the linear matrix space, ensuring positive semi-definiteness, which avoids geometrical distortions. Mathematically, this Jacobian Blend Skinning is defined as:
$\mathbf{J}_{b}:=\text{JBS}(\mathbf{J},w)=\underbrace{\text{exp}(\sum_{i\in adj.}w_{i}\text{log}(\mathbf{U}_{i}))}_{\mathbf{U}_{b}}\cdot\underbrace{\sum_{i\in adj.}w_{i}\mathbf{P}_{i}}_{\mathbf{P}_{b}}.$ (2)
@@
The blending weights, $w_{i}$, in JBS are learnable convex weights that control how much influence each adjacent triangle has on the resulting blended Jacobian $\mathbf{J}_{b}$. Note that since $w_{i}$ is activated by Sigmoid function, the blended stretch/shear component $\mathbf{P}_{i}$ is also guaranteed positive semi-definite property. The exponential and logarithm mappings for rotations are performed using Rodrigues’ formula. Finally, the blended Jacobian $\mathbf{J}_{b}$ is then used to deform the canonical Gaussian’s covariance and mean position and also transform the normal.
@@ 4.2 Head Avatar Reconstruction and Reenactment
Table 1: Qualitative comparison on FaceTalk (Zheng et al., 2022).
@@ 4.3 Ablation Study
We further validate the effectiveness of our method’s Jacobian and Jacobian Blend Skinning (JBS) in Tab. 3 and Fig. 6(a). Here,“Vanilla” refers to the combination of 2DGS (Huang et al., 2024) and Gaussian Avatars (Qian et al., 2024), with eyeball convexity regularization and ASGs as described in Sec. 2.4.
@@ A.1 Related Work
Reconstructing and Animating Head Avatars. Existing approaches for reconstructing and animating avatars mainly differ in two fundamental aspects: implicit or explicit models. Implicit models reconstruct the face by neural radiance field in combination with volumetric rendering or using implicit surface functions. (Gao et al., 2022), (Zielonka et al., 2023b), (Zheng et al., 2022), (Xu et al., 2023) With explicit models (Grassal et al., 2022),(Zheng et al., 2023),(Khakhulin et al., 2022), the seminal work of 3DMFM uses principal component analysis (PCA) to model facial appearance and geometry on a low-dimensional linear subspace. 3DMFM and its variants have been widely applied in optimization-based and deep learning-based head avatar creation. Recently, due to its efficient rendering and topological flexibility, there are many works utilizing 3D Gaussian Splatting (Kerbl et al., 2023). These works can be further categorized based on how they define and use Gaussians. Some approaches bind Gaussians directly to the mesh (Shao et al., 2024; Qian et al., 2024), while others map them onto UV coordinates (Xiang et al., 2024). In some cases, Gaussians are extracted as features (Giebenhain et al., 2024),(Xu et al., 2024), or a neural parametric model replaces the traditional 3DMFM (Giebenhain et al., 2024). However, in most cases, the deformation is often handled by the mesh itself with little additional consideration.
@@ A.3 Implementation Details
A.4 Additional Ablation Study
@@
A.5 Limitation and Seminal Future Work
@@
Second, the inherent limitations of 3DMFM’s representation, particularly the absence of hair strands and the tongue, can degrade the render quality. Since 3DMFM is fundamentally a bald model for human heads, it lacks the capability to represent the high-frequency details of human hair (Fig.  10). We believe that integrating our model with parametric strand models (Sklyarova et al., 2023; Zakharov et al., 2024; Rosu et al., 2022) could help alleviate these issues.
@@
A.6 Conceptual comparison with previous method

### v2

@@ 2.1.1 Gaussian Splatting
3D Gaussian Splatting (Kerbl et al., 2023) (3DGS) reconstructs a scene with anisotropic 3D Gaussian primitives. Each Gaussian is defined by a positive semi-definite covariance matrix $\Sigma$ that is centered at a position $\mu$. The covariance $\Sigma$ is decomposed as $\Sigma=RSS^{T}R^{T}$, where $R$ is a rotation matrix and $S$ is a diagonal scaling matrix. This covariance matrix is used when estimating each Gaussian’s function value at 3D point $x$ such as $G(x)=\text{exp}({-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)})$. Besides these geometrical properties, each Gaussian has appearance properties opacity $\alpha$ and color $c$. To render the scene, the final color $C$ is computed by alpha-blending Gaussians after projecting them to the image plane: $C=\sum_{i=1}c_{i}\alpha_{i}G^{proj}(x)\prod_{j=1}^{i-1}(1-\alpha_{i}G^{proj}(x)).$ $G^{proj}$ is given by evaluating the function value of 2D projection of the 3D Gaussian in image space by EWA volume splatting algorithm (Zwicker et al., 2001).
@@ 2.2 Affine transformation of 2D surfels
Local Geometry Descriptor. Let us consider a triangle in the canonical space as the matrix formed by its vertices ${v_{i}}\in\mathbb{R}^{3}$, and its edge matrix as $\mathbf{E}=[v_{1}-v_{0},v_{2}-v_{0},v_{3}-v_{0}]\in\mathbb{R}^{3\times 3}$, where $v_{3}:=v_{0}+\frac{(v_{1}-v_{0})\times(v_{2}-v_{0})}{\sqrt{|(v_{1}-v_{0})\times(v_{2}-v_{0})|}}$. $\tilde{\mathbf{E}}$ is the deformed version with the deformed triangle’s vertices $\tilde{v_{i}}$ following Eq. 1. Note that in GaussianAvatars, the edge matrix is defined with a normalized base, height, and normal direction. This constructs an orthogonal matrix and approximates the shear or stretching-related deformations with isotropic scaling from the average of the base and height length. Please find a detailed derivation in GaussianAvatars (Qian et al., 2024).
@@ 2.3 Jacobian Blend Skinning (JBS)
Replacing the scaled-rotation component $s_{p}R_{p}$ of the similarity transform with the proposed $\mathbf{J}$ enables better representation of shear and stretch. However, ensuring the continuity of adjacent deformations is crucial for achieving smoother transformations. To address this, we introduce an improved technique to mitigate such discontinuities by substituting $\mathbf{J}$ again with a blended Jacobian, $\mathbf{J}_{b}$. We refer to this approach as Jacobian Blend Skinning (JBS), which builds on the principles of Linear Blend Skinning (LBS) (Badler & Morris, 1982) while addressing its inherent limitations.
@@
Degenerate Solution of Linear Blend Skinning in Matrix Space. Linear Blend Skinning (LBS) is widely used for dynamic modeling of the human body and head (Li et al., 2017; Loper et al., 2015), but it struggles with the non-linearity of rotation matrix interpolation in $SO(3)$, causing distortions in rotational transformations. Our toy experiments (Fig. 2(b)) show how element-wise linear interpolation distorts shapes. Previous works (Zielonka et al., 2023b; Shao et al., 2024) also adopt this sub-optimal approach to handle local deformations. Techniques like matrix-logarithm interpolation or quaternion’s SLERP can address this but are limited to rotations.
@@
Introducing Jacobian Blend Skinning (JBS). The Jacobian Blend Skinning (JBS) algorithm overcomes the limitations of LBS by focusing on the interpolation of Jacobian gradients, which encapsulate not only rotations but also shear and stretch, both residing in the broader $GL(3)$ (General Linear) field. The key to JBS is leveraging Polar Decomposition (PD) (Shoemake & Duff, 1992) to break down Jacobian gradient $\mathbf{J}$ into two meaningful components: rotation (orthogonal matrix $\mathbf{U}$) and stretch/shear (symmetric positive semi-definite matrix $\mathbf{P}$): $\mathbf{J}=\mathbf{U}\mathbf{P}$. This decomposition is unique and coordinate-independent (proof in Appendix A.2.1), making it geometrically sound for interpolation. By blending the rotation in matrix-logarithm space and the stretch/shear in linear matrix space, JBS ensures that the resulting transformations remain geometrically valid, avoiding the distortions seen with LBS (Fig. 2(b)).
@@
JBS is applied to the Gaussian deformations considering not only its parents’ triangle, but its adjacent triangles. The blending itself is performed separately for the rotational ($\mathbf{U}_{b}$) and stretch/shear ($\mathbf{P}_{b}$) components. For rotations, the blending weights are applied in the matrix-logarithm space $\mathfrak{so}(3)$, followed by an exponential mapping back into the $SO(3)$ space. For stretch/shear, the weights are applied directly in the linear matrix space, ensuring positive semi-definiteness, which avoids geometrical distortions. Mathematically, this Jacobian Blend Skinning is defined as:
$\mathbf{J}_{b}:=\text{JBS}(\mathbf{J},w)=\underbrace{\text{exp}(\sum_{i\in adj.}w_{i}\text{log}(\mathbf{U}_{i}))}_{\mathbf{U}_{b}}\cdot\underbrace{\sum_{i\in adj.}w_{i}\mathbf{P}_{i}}_{\mathbf{P}_{b}},$ (2)
@@
where $adj.$ implies the set of adjacent triangles. The blending weights, $w_{i}$, in JBS are learnable convex weights that control how much influence each adjacent triangle has on the resulting blended Jacobian $\mathbf{J}_{b}$. Note that since $w_{i}$ is activated by Sigmoid function, the blended stretch/shear component $\mathbf{P}_{i}$ is also guaranteed positive semi-definite property. The exponential and logarithm mappings for rotations are performed using Rodrigues’ formula. Finally, the blended Jacobian $\mathbf{J}_{b}$ is then used to deform the canonical Gaussian’s covariance and mean position and also transform the normal.
@@ 4.2 Head Avatar Reconstruction and Reenactment
Table 1: Quantitative comparison on FaceTalk (Zheng et al., 2022).
@@ 4.3 Ablation Study
We further validate the effectiveness of our method’s Jacobian and Jacobian Blend Skinning (JBS) in Tab. 3 and Fig. 6(a). Here,“Vanilla” refers to the combination of 2DGS (Huang et al., 2024) and Gaussian Avatars (Qian et al., 2024).
@@ 5 Conclusion and Discussion
Acknowledgment
This research was supported by the Institute for Information & Communications Technology Planning & Evaluation (IITP) grant, funded by the Korean government (MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program at KAIST). We also acknowledge the support of the National Research Foundation of Korea (NRF) grant, funded by the Korean government (MSIT) (No. RS-2025-00555621). Additionally, we would like to express our sincere gratitude to Jungwoo Ahn for conducting the baseline experiments.
Reproducibility statement
To ensure the reproducibility of our method, we provide the anonymous GitHub link. Moreover, the details on implementation and the information of datasets can be found in the appendix A.3 and Sec. 4.1.2, respectively. For NeRSemble (Kirschstein et al., 2023) dataset, we used pre-processed dataset from GaussianAvatars (Qian et al., 2024). Together with provided code and publicly available datasets, we believe our paper contains sufficient information for reimplementation.
Ethics statement
The creation of highly accurate and geometrically detailed head avatars raises significant ethical and security concerns, particularly regarding privacy violations. The realistic nature of these avatars can lead to unauthorized manipulation, where personal likenesses may be misused without consent. This includes altering appearances in virtual environments or generating misleading content, which infringes on privacy and may result in serious consequences.
One critical issue is the potential for creating deepfake videos, which could spread false information, harm reputations, and manipulate public opinion. The increased realism of these avatars makes it difficult to discern authentic content from fabricated material, amplifying risks of misinformation and defamation. Additionally, the possibility of identity theft through the exploitation of digital avatars presents a serious threat to personal security and public trust in digital media.
To mitigate these risks, we will strictly limit the use of our method to academic research. This restriction aims to prevent misuse and ensure the technology is applied in an ethical and constructive manner.
@@ A.1 Related Work
Neural Surface Reconstruction. Emerging from the success of NeRF (Mildenhall et al., 2021), neural surface reconstruction (Oechsle et al., 2021; Yariv et al., 2020; Wang et al., 2021) has garnered significant attention from researchers, leveraging Occupancy Fields and Signed-Distance Functions (SDFs). Notably, subsequent works (Rosu & Behnke, 2023; Li et al., 2023) have introduced methodologies that integrate fast and efficient hash-encoding-based radiance fields (Müller et al., 2022) with SDFs. This development has inspired research into human head geometry reconstruction, moving beyond mesh-based topology-invariant 3D Morphable Face Models (3DMFMs) (Li et al., 2017; Paysan et al., 2009). The first major work in this area, H3D-Net (Ramon et al., 2021), proposed an SDF-based approach for few-shot human head geometry reconstruction using an auto-decoder-trained model. Building on this, a subsequent study (Xu et al., 2023a) targeted a similar setting as H3D-Net, focusing on more detailed human head geometry reconstruction with single-person refinement. Recently, MonoNPHM (Giebenhain et al., 2024a) introduced a few-shot head reconstruction framework based on a pretrained neural parametric head model (Giebenhain et al., 2023), which combines the topology-free advantages of SDFs with the manipulability of parametric models. Departing from these SDF-based approaches, we are the first to propose reconstructing dynamic head geometry from RGB videos, including monocular setups, utilizing Gaussian primitives (Huang et al., 2024).
Reconstructing and Animating Head Avatars. Existing approaches for reconstructing and animating avatars mainly differ in two fundamental aspects: implicit or explicit models. Implicit models reconstruct the face by neural radiance field in combination with volumetric rendering or using implicit surface functions. (Gao et al., 2022), (Zielonka et al., 2023b), (Zheng et al., 2022), (Xu et al., 2023b) With explicit models (Grassal et al., 2022),(Zheng et al., 2023),(Khakhulin et al., 2022), the seminal work of 3DMFM uses principal component analysis (PCA) to model facial appearance and geometry on a low-dimensional linear subspace. 3DMFM and its variants have been widely applied in optimization-based and deep learning-based head avatar creation. Recently, due to its efficient rendering and topological flexibility, there are many works utilizing 3D Gaussian Splatting (Kerbl et al., 2023). These works can be further categorized based on how they define and use Gaussians. Some approaches bind Gaussians directly to the mesh (Shao et al., 2024; Qian et al., 2024), while others map them onto UV coordinates (Xiang et al., 2024). In some cases, Gaussians are extracted as features (Giebenhain et al., 2024b),(Xu et al., 2024), or a neural parametric model replaces the traditional 3DMFM (Giebenhain et al., 2024b). However, in most cases, the deformation is often handled by the mesh itself with little additional consideration.
@@ A.3 Implementation Details
A.4 Additional Experiement
Comparsion with Gaussian Head Avatars (GHA) (Xu et al., 2024) GHA serves as a strong baseline in rendering quality, leveraging an additional super-resolution model in screen space. As shown in Tab. 6, GHA outperforms our method in terms of PSNR. However, in other metrics such as SSIM and LPIPS, GHA falls behind.
Table 6: Quantitative comparsion with GHA (Xu et al., 2024)
[…]
GHA 27.25 0.909 0.153 0.505
Ours 26.20 0.932 0.052 0.837
Fig. 8 reveals potential reasons for this discrepancy. Notably, artifacts such as over-saturation are visible, which we attribute to GHA’s screen-space refinement. This refinement struggles when rendering extreme poses that fall outside the distribution expected by the super-resolution model. Additionally, GHA faces challenges in reconstructing high-fidelity geometry and handling extreme expressions, both of which are key strengths of our approach. We would like to emphasize the importance of capturing the specular highlights of the eyes, particularly the cornea. GHA renders the pupils with a matte appearance, neglecting the high-frequency specular reflections in the corneal region. We argue that these specular details are critical for enhancing the realism of the eyes, which play a pivotal role in creating immersive and lifelike head avatars.
Figure 8: Qualitative Comparison with GHA (Xu et al., 2024)
A.5 Additional Ablation Study
@@
Training and Rendering Time Complexity Measurement Tab. 7 summarizes the rendering speed and training time for our method and GaussianAvatars (Kerbl et al., 2023). The Base configuration refers to replacing the 3D Gaussian Splatting rasterizers in GaussianAvatars with their 2D counterparts. Our method incurs only a 17% drop in rendering speed and an additional 25 minutes of training time, while still achieving 3$\times$ real-time rendering speeds (generally over 30 FPS) and maintaining efficient training. A detailed analysis attributes the minimal training and testing overhead to our GPU-level CUDA kernel implementation for Jacobian computations.
For rendering, the primary factor behind the speed reduction is the Jacobian Blend Skinning (JBS), where the overhead mainly arises from the Polar Decomposition step. Our current implementation utilizes PyTorch’s SVD, which relies on the cuSOLVER backend. To further investigate this bottleneck, we conducted additional experiments using RoMA (Brégier, 2021)’s specialized Procrustes routine, which is designed to efficiently compute the $3\times 3$ unitary matrix $\mathbf{U}$ of the Jacobian $\mathbf{J}$. Notably, replacing torch.svd with roma.special-procrustes yielded a performance gain of approximately 4–5 FPS.
Although this demonstrates the potential of alternative approaches, there is still room for further improvement. Higham’s routine (Higham & Noferini, 2016), specifically tailored for $3\times 3$ matrices, offers a promising direction to address this overhead and is well-suited for CUDA-based implementations.
Method GA Base + Jacobian +JBS +SGs (= Ours)
Rendering Speed (FPS) +PyTorch 71.18 109.36 107.59 92.62 90.13
+RoMA N/A N/A N/A 97.29 94.72
Training Time (hours) 1.65 1.68 1.73 1.98 2.11
Table 7: Comparison of training and rendering (test) computational costs.
A.6 Limitation and Seminal Future Work
@@
Second, the inherent limitations of 3DMFM’s representation, particularly the absence of hair strands and the tongue, can degrade the render quality. Since 3DMFM is fundamentally a bald model for human heads, it lacks the capability to represent the high-frequency details of human hair (Fig.  11). We believe that integrating our model with parametric strand models (Sklyarova et al., 2023; Zakharov et al., 2024; Rosu et al., 2022) could help alleviate these issues.
@@
A.7 Conceptual comparison with previous method